OpenStudy (mysticfury):

What is the solution to the system of equations represented by the two equations? y = 3/2x y = -1/2x + 4 (__ , __)

OpenStudy (mysticfury):

I have no clue how to do this Substitution stuff

OpenStudy (shaik0124):

substitute eqn 1 y in eqn 2 u will get x then putx in any one equation u will get y

OpenStudy (mysticfury):

idk how to do that

OpenStudy (shaik0124):

3/2 x =-1/2x+4 solve for x

OpenStudy (shaik0124):

multiply 2 on both sides now

OpenStudy (mysticfury):

umm do i add -1/2 and 4? then multiply by 2?

OpenStudy (shaik0124):

no multiply by 2 first then iu will get 3x=-1x+4

OpenStudy (shaik0124):

sorry 3x=-1x+8

OpenStudy (mysticfury):

oh okay i was confused what number on the left side to multiply by 2

OpenStudy (mysticfury):

right side sorry

OpenStudy (shaik0124):

x=2

OpenStudy (mysticfury):

how did you get that?

OpenStudy (shaik0124):

subst x=2 in any eqn u will get y

OpenStudy (shaik0124):

3x+1x=-1x+1x+8

OpenStudy (mysticfury):

so y is 3?

OpenStudy (shaik0124):

added 1x on both sides

OpenStudy (shaik0124):

yeah

OpenStudy (mysticfury):

so my answer is 2/3? thats the same as the last question >.>

OpenStudy (mysticfury):

thank you @shaik0124

OpenStudy (shaik0124):

yes 2/3

OpenStudy (shaik0124):

welcome